Akira earns a salary of $200 a week plus 15% comission on all sales. Juanita, new to the company, has a straight commision salary of 25%.
A. Model each salary with a linear equation. Graph both equations on the coordinate plane.
B. Last week, Akira and Juanita had the same amount of sales but Juanita's salary was $100 more than Akira's, what was the amount of sales?
C. Suppose Hector, also new to the company, is offered a straight commission of 15% can he ever make as much as Akira? Justify your answer.
Please Explain all answers! I really would like to learn it. Thank You!

Answer by stanbon(48558)

Akira earns a salary of $200 a week plus 15% comission on all sales. Juanita, new to the company, has a straight commision salary of 25%.
A. Model each salary with a linear equation. Graph both equations on the coordinate plane.
Akira Salary: S(x) = 200 + 0.15x, where x is her total sales.
Juanita Salary: J(x) = 0.25x, where x is her total sales
--------------------------------------------------------------------
B. Last week, Akira and Juanita had the same amount of sales but Juanita's salary was $100 more than Akira's, what was the amount of sales?
Juanita Salary - Akira Salary = 100
0.25x - (200+0.15x) = 100
0.10x - 200 = 100
0.10x = 300
x = $3000 (amount of sales)
-------------------------------------
C. Suppose Hector, also new to the company, is offered a straight commission of 15% can he ever make as much as Akira? Justify your answer.
Hector's Salary: H(x) = 0.15x
Equation:
Hector Salary = Akira Salary
0.15x = 200 + 0.15x
Akira's Salary will always be $200 more than Hector's.
Cheers,
Stan H.